Claude and Rosalinda are debating whether an equation from their homework assignment is an indentily .Claude says that some he has tried ten specific value for the variable and all of them worked ,it must be used as counterexamples Roselinda explain that specific values could only be use as counterexamples to prove that an equation is not an identity.Who is correct? .Explain your answer.
Answer by Edwin McCravy(20056)

Rosalinda was the one who is right. Here is an example of an equation 

(x-1)(x-2)(x-3)(x-4)(x-5)(x-6)(x-7)(x-8)(x-9)(x-10) = 0


that holds true when you substitute

x=1, x=2, x=3, x=4, x=5, x=6, x=7, x=8, x=9, and x=10,

for you get 0 = 0 with all 10 of those values.
However when you substitute x = 11 in it, you get

            3628800 = 0

So it's not an identity even though it works for all those 10 values of x.  To be an identity it
must work for all values of x.  x=11 is a
counterexample which proves it is not an identity.

An identity is like this:

   2x + 3 = 3(x + 1)-x 

Holds for ALL values of x, not just ten.
You cannot find a counterexample like you
can with the first one above.

Edwin